A brand of grape drink that is 25% grape juice is mixed with another brand that is 75% grape juice. The result is 7 gallons of grape drink that is 35% grape juice. How much of the brand that is 25% grape juice was used? What are the steps to determine the answer?
Answer by stanbon(75887)

A brand of grape drink that is 25% grape juice is mixed with another brand that is 75% grape juice. The result is 7 gallons of grape drink that is 35% grape juice. How much of the brand that is 25% grape juice was used? What are the steps to determine the answer?
--------------
Comment: grape drink is made up of grape juice and water. You need to
keep track of the grape juice or the water in each of the solutions.
--------------------
25% grape juice DATA:
Amount = x gallons ; Amt of grape in it is 0.25x gallons
-----------------
75% grape juice DATA:
Amount = "7-x" gallons; Amt of grape in it is 0.75(7-x) = 5.25-0.75x gallons
-------------------
35% mixture DATA:
Amount = 7 gallons: amt of grape is 0.35*7 = 2.45 gallons
-------------------
EQUATION:
grape + grape = grape
0.25x + 5.25-0.75x = 2.45
-0.50x = -2.8
x = 5.6 gallons (amount of 25% grape juice that must be added)
================
Cheers,
Stan H.